# Model Predictive Control via Probabilistic Inference: A Tutorial

### v1

@@ 4 Applications and Extensions
This equation indicates that the optimal control distribution is expressed as the product of the Boltzmann distribution and the prior distribution (Eq. (2)), and that the variational distribution approximates this optimal control distribution by minimizing the KL divergence. We can generate control input sequences by sampling from the variational distribution via Eq. (9). This section discusses key points for improving performance and tuning, as well as advanced research topics based on this general framework.
4.1 Keys to Performance Improvement and Tuning
Revisiting Eq. (14), for a given specific task and robot, the trajectory cost $J_{\tau}$ and the control input $\mathbf{u}_{0:T-1}$ are uniquely determined. Additionally, the optimal control distribution is theoretically given by Eq. (2), although there have been discussions regarding the shape of the optimality likelihood.
Therefore, the design elements we need to consider are the temperature parameter $\lambda$, the prior distribution $p(\mathbf{u}{0:T-1})$, and the variational distribution $\pi{\theta}$. Among these, as discussed in Section 3.2.3, the temperature parameter is a hyperparameter that governs the trade-off between stochastic and deterministic behavior. In contrast, the design of the prior and variational distributions has a significant influence on control performance and characteristics. Based on this insight, we discuss these two design aspects, i.e., the prior and variational distributions, in detail below.
4.1.1 Prior Distribution $p(\mathbf{u}_{0:T-1})$ is “Search Region for Solutions”
Figure 7: Variation of the optimal control distribution with the prior distribution. The cost function is the same as in Fig. 4. Four cases of the prior distribution $p(u)=\mathcal{N}(\mu,\sigma)$ are shown with $\mu=\{-3.0,0.0\}$ and $\sigma=\{0.25,1.0\}$.
As discussed in Section 3.2.1, the optimal control distribution is regularized by the prior distribution, making the prior a crucial element determining the optimal control distribution’s shape.
Figure 7 shows how the optimal control distribution changes with variations in the prior distribution’s mean and variance. Increasing the deviation between the prior distribution and the Boltzmann distribution’s principal mode causes the optimal control distribution to deviate from the Boltzmann distribution, degrading optimality in the sense of cost function minimization (bottom-right and top-right panels). Additionally, increasing the prior distribution’s variance moves the optimal control distribution closer to the Boltzmann distribution when their deviation is large (top-right), which is desirable. However, when their deviation is small (top-left), the resulting optimal control distribution also has large variance, lowering the probability of generating the optimal control distribution’s principal mode and potentially failing to achieve optimal actions in the cost function sense. Thus, the prior distribution to avoid most is “one with large deviation from the Boltzmann distribution and small variance” (bottom-right), while the most desirable prior is “one close to the Boltzmann distribution with small variance” (bottom-left).
This becomes more intuitive from an algorithmic perspective. As shown in Eq. (12), the prior distribution defines the “search region for solutions” from which we actually sample. Therefore, it is desirable to focus sampling around the vicinity of the Boltzmann distribution’s mode peaks.
In summary, the prior distribution must be designed with awareness of the Boltzmann distribution (i.e., the cost function). Consequently, various approaches have been proposed, including cost-aware adaptive priors [honda2024stein, yi2024covo], optimization via learning [power2022variational], and utilization of reinforcement learning policies [hansen2022temporal].
4.1.2 Variational Distribution $\pi_{\theta}$ is “Diversity of Behaviors”
The variational distribution approximates the optimal control distribution and ultimately generates control inputs. Therefore, the variational distribution determines the “diversity of generated actions”. For example, MPPI assumes a unimodal Gaussian variational distribution, which cannot capture multiple modes of the behavior, i.e., a multimodal optimal control distribution, generating solutions only near a single mode.
To improve behavior diversity, more flexible variational distributions are needed. Consequently, Gaussian mixture distributions [okada2020variational] and particle-based representations [lambert2021stein] have been proposed as variational distributions. These methods enhance exploration capabilities through increased solution diversity, facilitating model-based reinforcement learning [okada2020variational]. However, they tend to suffer from increased computational cost and reduced stability.
4.2 Advanced Research Topics
4.2.1 Theoretical Analysis
Probabilistic inference MPC is even more challenging to analyze theoretically than numerical optimization-based MPC. However, recent years have seen a small number of theoretical analysis studies. Reference [yi2024covo] proves that MPPI guarantees linear convergence when the number of samples is infinite and the cost function is quadratic. However, infinite samples are impractical. It is important to note that with finite samples, convergence is not guaranteed even for quadratic cost functions due to sample complexity effects (Section 3.2.4).
4.2.2 Accelerating Decision-Making
As discussed in Section 3.2.2, optimal control distributions in probabilistic inference MPC tend to exhibit delayed decision-making. Additionally, the variational distribution obtained in Eq. (9) minimizes the forward KL divergence. Forward KL divergence minimization corresponds to maximum likelihood estimation, causing the resulting variational distribution to tend to cover multiple modes. Consequently, decision-making is further delayed.
This decision-making delay can sometimes lead to dangerous behaviors. For instance, when controlling fast-moving vehicles, optimal behavior involves avoiding obstacles with minimal clearance, so even slight dynamics errors may cause collisions. To accelerate decision-making, methods that minimize reverse KL divergence instead of forward KL divergence in Eq. (9) have been proposed [honda2024stein, kobayashi2022real]. Reverse KL divergence minimization tends to cover only a single mode rather than multiple modes, enabling faster decision-making.
4.2.3 Connection to Diffusion Models
Recently, diffusion models that generate diverse data through denoising have attracted attention, and connections between MPPI and diffusion models have been identified [xue2024full]. Literature [xue2024full] demonstrates that when solving the optimal control distribution using score-based diffusion models, part of the score function can be replaced with MPPI computation, enabling sample generation by solving MPPI for the number of diffusion steps without requiring training. This method can capture smaller modes compared to MPPI and has been reported to achieve high performance even in high-dimensional control problems such as quadruped robots. However, the need to solve MPPI for each diffusion step results in very high computational cost per sample generation, which remains a challenge for future work.

### v2

@@ 2.6 Summary of the Tutorial
This relation highlights two key points. First, the optimal control distribution can be written as the product of a Boltzmann term and a prior (Eq. (2)). Second, we approximate this target distribution with a parameterized variational distribution by minimizing the KL divergence. Control inputs are then obtained by sampling from the resulting variational distribution as in Eq. (9).
For a given specific task and robot, the trajectory cost $J_{\tau}$ and the control input $\mathbf{u}_{0:T-1}$ are uniquely determined. Additionally, the optimal control distribution is theoretically given by Eq. (2). Therefore, the main design choices are the prior $p(\mathbf{u}_{0:T-1})$, and the variational family $\pi_{\theta}$. Based on this framework, we organize the challenges and methods in PI-MPC as follows.
3 Survey of Probabilistic Inference-Based MPC Methods
As established in Section 2, the main design choices in PI-MPC reduce to two elements: the prior distribution $p(\mathbf{u}_{0:T-1})$, and the variational family $\pi_{\boldsymbol{\theta}}$. Each choice involves distinct trade-offs between solution quality, computational cost, and robustness. This section surveys the key challenges that arise when instantiating these choices for real-world robotic tasks, together with representative methods that address them.
We organize the discussion as follows. Sections 3.13.3 address algorithmic design challenges: prior distribution design (Section 3.1), multi-modal distribution handling (Section 3.2), and constraint satisfaction (Section 3.3). Sections 3.43.5 turn to practical scalability, covering high-degree-of-freedom robots and hardware acceleration. Finally, Sections 3.6.23.7 examine theoretical foundations and connections to related paradigms.
3.1 Design of Prior Distributions
Figure 7: Effect of prior distribution mismatch and variance on the optimal control distribution. Larger deviation between the prior and the Boltzmann distribution deteriorates optimality, while increased prior variance mitigates this effect when mismatch is large. Priors located near the Boltzmann mode with small variance yield the most favorable performance, whereas distant low-variance priors produce the poorest outcome. The cost function is the same as in Fig. 4, and four Gaussian prior distributions $p(u)=\mathcal{N}(\mu,\sigma),\mu\in\{-3.0,0.0\}$ and $\sigma\in\{0.25,1.0\}$ are compared.
The prior distribution $p(\mathbf{u}_{0:T-1})$ serves a dual role in PI-MPC: it defines the sampling region from which candidate solutions are drawn, and it acts as a regularizer that shapes the notion of optimality through the KL penalty in Eq. (5). A poorly chosen prior wastes samples in low-probability regions of the optimal distribution, while an overly narrow prior risks missing the global optimum entirely. The principal lever for addressing this trade-off is to bring the prior close to the modes of the Boltzmann distribution while keeping its variance small enough for efficient sampling yet large enough to avoid premature convergence.
Existing methods can be grouped into four categories: (i) shaping the prior for a better exploration–exploitation balance, (ii) guiding samples with auxiliary controllers, (iii) adapting the prior online in a cost-aware manner, and (iv) learning the prior from data.
3.1.1 What is a Good Prior Distribution?
Before reviewing specific methods, we establish the design objective using a concrete example. Figure 7 illustrates how the optimal control distribution varies with the mean and variance of the prior. When the prior mean is far from the principal mode of the Boltzmann distribution, the resulting optimal control distribution shifts away from low-cost regions, degrading optimality (bottom-right and top-right panels). Increasing the prior variance partially compensates for this mismatch by spreading samples more broadly (top-right), but when the prior is already well-centered, a large variance dilutes probability mass around the mode and reduces the chance of generating near-optimal actions (top-left). The worst case is a prior that is both far from the Boltzmann mode and tightly concentrated (bottom-right); the best case is a prior centered on the mode with small variance (bottom-left).
From an algorithmic standpoint, this picture is intuitive: in the MPPI algorithm (Section 2.5), the prior defines the search region from which all samples are drawn. Focusing this region around the modes of the Boltzmann distribution maximizes both sample efficiency and solution quality, which is supported by theoretical analyses [Yoon2022-jb, Tao2022-dj]. The methods reviewed below all aim, in different ways, to achieve this ideal: a prior with small variance positioned near an optimal mode (Fig. 7, bottom-left).
3.1.2 Balancing Exploration and Exploitation via Prior Distribution
Reducing the variance of a standard Gaussian prior improves sample efficiency and solution quality, but it also narrows the search region. If the prior is misaligned with the true mode, a small variance can trap in a suboptimal region and slow convergence. The key idea is therefore to reshape the prior so that its variance is small in a meaningful sense, i.e., concentrating samples around promising regions, while retaining sufficiently heavy tails to avoid missing distant modes.
One family of methods achieves this by replacing the Gaussian prior with a leptokurtic distribution, i.e., one with higher kurtosis and heavier tails. Mohamed et al.  [mohamed2022autonomous] used a normal–log-normal (NLN) mixture, and Vlahov et al.  [vlahov2024low] employed colored noise as the prior. Such distributions suppress high-frequency components while preserving low-frequency structure, yielding temporally smooth samples that concentrate around plausible trajectories without sacrificing coverage.
Temporal smoothness can also be imposed more directly. Post-processing filters [andrejev2025pi] and spline interpolation [bhardwaj2022storm] smooth samples after generation, while other methods redefine the action space in terms of action derivatives [kim2022smooth, lee2time], so that sampling in the new space inherently produces smooth trajectories. A more principled alternative is to embed temporal correlations into the inference problem itself: Gaussian Processes (GPs) naturally encode such correlations and can replace the temporally independent Gaussian prior used in standard MPPI [mukadam2018continuous, urain2022learning, watson2023inferring].
3.1.3 Guiding Samples with Auxiliary Controllers
Rather than reshaping the prior itself, another strategy is to steer samples toward promising regions using feedback from auxiliary controllers. Each sampled trajectory is corrected by a lightweight controller that pushes it toward a mode of the optimal control distribution, effectively narrowing the sampling region without reducing the nominal prior variance.
When linear approximations of dynamics and cost are available, analytic feedback controllers offer a computationally cheap solution. Examples include task-specific Linear Quadratic Regulators (LQR) [trevisan2024biased], iterative Linear Quadratic Gaussian (iLQG) controllers [williams2018robust, gandhi2021robust], and covariance steering controllers [yin2022trajectory]. For problems where such approximations are unavailable, gradient-based methods provide a more general alternative: convex optimization [rastgar2024priest], Stochastic Gradient Descent (SGD) [bharadhwaj2020model], and Adam [heetmeyer2023rpgd] can guide samples using only gradient information of the cost and dynamics. For the most general setting, i.e., where even cost and dynamics gradients may not be available, Stein Variational Gradient Descent (SVGD) offers a derivative-free option that approximates the KL gradient from samples alone [lambert2021stein, honda2024stein, miura2024spline].
3.1.4 Cost-Aware Adaptive Prior Distribution
Instead of correcting individual samples, a complementary strategy is to adapt the prior distribution itself in response to observed trajectory costs. Because the adaptation modifies the distribution rather than individual samples, its computational cost is independent of the sample count, and the resulting algorithm remains consistent with the PI-MPC theoretical framework.
The central difficulty is a circular dependency: evaluating the quality of a prior requires drawing i.i.d. samples and rolling them out to compute trajectory costs, but these costs are themselves influenced by the prior. Asmar et al.  [asmar2023model] break this circularity through iterative refinement, i.e., alternating between cost evaluation and prior update, at the expense of higher computational cost. Lighter-weight alternatives avoid full iteration: Tao et al.  [tao2022path] use Stochastic Control Barrier Functions (SCBF) to set a safe mean and variance, while Honda et al.  [honda2024stein] employ SVGD on a small set of pilot samples to capture the mode shape before the main sampling phase. Other approaches sidestep the circularity by predicting future cost characteristics: Mohamed et al.  [mohamed2025towards] propagate variance dynamics via the unscented transform and fold the result into the cost, and Yi et al.  [yi2024covo] derive a closed-form optimal covariance under simplifying assumptions on the cost structure.
3.1.5 Learning-based Prior Distribution
When offline data or simulation is available, machine learning can be used to construct priors that are already well-aligned with the task before online execution begins.
If expert demonstrations are available, generative models trained on the dataset can serve directly as priors. Urain et al.  [urain2022learning] learn an energy-based model of the optimal trajectory distribution, while Carvalho et al.  [carvalho2023motion] use diffusion models [ho2020denoising] to capture richer multi-modal structure. A complementary line of work incorporates model-free RL policies as priors: pre-trained policies can initialize the MPPI mean [qu2023rl, wang2025residualmppi], or MPPI can be embedded as a residual correction during RL training so that the learned policy and the MPC prior co-evolve [sacks2024deep]. In model-based RL, MPPI serves as the planner while a model-free policy provides the prior [hansen2022temporal].
More specialized approaches learn environment-specific priors using Normalizing Flows [power2022variational, sacks2023learning]. By incorporating the KL objective in Eq. (4) into the training loss, these methods can learn priors in a self-supervised manner without expert data, conditioning only on environment features and start–goal pairs.
3.2 Handling of Multi-Modal Distributions
Multi-modality of the optimal control distribution arises naturally in robotics whenever multiple solutions have comparable costs: an autonomous vehicle can pass an obstacle on either side and a manipulator can choose among several grasp poses. Failing to handle multi-modality can produce dangerous mode-averaging behavior. For example, steering directly into an obstacle rather than clearly committing to one side. The principal lever is the choice of variational family $\pi_{\boldsymbol{\theta}}$ and the divergence used to fit it. Existing approaches fall into two categories: (i) approximating the full multi-modal distribution with an expressive variational family, and (ii) deliberately seeking a single mode for decisive action selection. The former is preferable when exploration and diversity are valuable (e.g., during RL), while the latter suits deployment scenarios that demand immediate, unambiguous decisions.
3.2.1 Variational Inference with Multi-modal Variational Distributions
The most direct strategy is to replace the unimodal Gaussian of MPPI with a more expressive variational family. Osa et al.  [osa2020multimodal] formulated offline trajectory optimization as cost-weighted density estimation using Gaussian Mixture Models (GMMs), and Okada et al.  [okada2020variational] adopted GMMs as the variational distribution within the PI-MPC framework, optimizing via mirror gradient descent. An alternative to parametric mixtures is to represent the variational distribution nonparametrically as a set of particles updated via SVGD [lambert2021stein, barcelos2021dual]. Wang et al.  [wang2021variational] generalized these particle-based methods by replacing the KL divergence with the Tsallis divergence, providing additional flexibility in controlling mode-covering behavior. Outside the strict VI framework, Carvalho et al.  [carvalho2023motion] integrated diffusion models into PI-MPC to generate multi-modal trajectory samples.
The primary strength of these methods is their ability to maintain solution diversity, which enhances exploration and can improve learning efficiency in model-based RL [okada2020variational]. Their main weakness relative to MPPI is computational: as shown in Eq. (12), MPPI obtains the globally optimal Gaussian approximation through a single round of sampling and weighted averaging, whereas multi-modal VI methods require iterative parameter updates, increasing both computational cost and the risk of numerical instability [^7] MPPI also requires iterative updates when simultaneously optimizing the mean and covariance [wang2021variational]..
3.2.2 Seeking a Single Mode in Multi-modal Distributions
When the goal is to execute a single action sequence, approximating the entire multi-modal distribution may be unnecessary. Instead, one can target a single high-quality mode and commit to it.
Clustering-based methods achieve this by partitioning samples post hoc. Park et al.  [park2025csc] employed DBSCAN to cluster samples by density, computed a weighted average within each cluster, and selected the cluster representative with the lowest cost. Zhang et al.  [zhang2024multi] computed separate weighted averages for multiple cost components and then re-weighted all samples to fuse them into a single control sequence.
A more principled alternative is to change the divergence objective. Standard PI-MPC minimizes the forward KL divergence $D_{\mathrm{KL}}(\pi^{*}\|\pi_{\boldsymbol{\theta}})$, which encourages the variational distribution to cover all modes (mode-covering). Minimizing the reverse KL divergence $D_{\mathrm{KL}}(\pi_{\boldsymbol{\theta}}\|\pi^{*})$ instead yields a mode-seeking solution that concentrates on a single mode. Because the reverse KL does not admit a closed-form solution, iterative methods such as SVGD [honda2024stein] and mirror descent [kobayashi2022real] are used in practice.
As discussed in Section 2.4.2, symmetry breaking causes the number of modes to decrease over time, so even mode-covering methods like MPPI eventually converge to a single mode. However, in dynamic environments or time-critical scenarios, this natural convergence may be too slow, and explicit mode-seeking can prevent dangerous hesitation.
3.3 Handling of Constraints
Real-world optimal control problems invariably involve constraints on both states and control inputs. Control input constraints (e.g., actuator limits) are relatively easy to enforce: samples that exceed bounds can be clipped after drawing from the prior [williams2018information], or inputs can be mapped through saturating functions such as $\tanh$; either approach can be absorbed into the dynamics of Eq. (1b). State constraints (e.g., collision avoidance, joint limits) are fundamentally harder because they couple the constraint with the forward dynamics and cannot be enforced by modifying a single sample in isolation. The simplest remedy is to add penalty terms to the cost function in Eq. (1a)—PI-MPC readily accommodates discontinuous penalties such as indicator functions and cost maps [macenski2020marathon2]—but penalties cannot guarantee strict satisfaction: if all samples from the prior violate a constraint, the “optimal” solution will itself be infeasible. Methods for stricter enforcement fall into two categories: (i) projecting samples onto the feasible set during prior sampling, and (ii) incorporating constraints directly into the variational inference problem. Projection methods are simpler to implement but may require differentiability or linearity assumptions; constrained VI is more principled but computationally heavier.
3.3.1 State-constrained Prior Sampling
The idea is to modify samples before cost evaluation so that they satisfy state constraints. In the MPPI framework, this typically means drawing Gaussian samples, rolling out the dynamics, and then projecting any constraint-violating trajectories onto the feasible set.
Several projection mechanisms have been proposed. Control Barrier Functions (CBFs) [gandhi2023safe, gandhi2022safety] project inputs to enforce forward-invariance of a safe set; primal–dual interior-point methods [park2025csc] and convex optimization [rastgar2024priest] solve small local programs to find the nearest feasible input. Yan et al.  [yan2024output] take the inverse approach: they sample directly in the state space and convert feasible state trajectories to control inputs via inverse dynamics.
For settings that require long-horizon safety guarantees, Borquez et al.  [borquez2025dualguard] project samples onto a future safety region defined by Hamilton–Jacobi reachability analysis, ensuring that the system can always return to a safe state. Wang et al.  [wang2025constrained] extend the framework to hierarchical equality and inequality constraints by combining the Lagrange multiplier method with null-space projection.
Projection-based methods are straightforward to implement and can enforce constraints strictly when the projection itself is exact. However, they often require differentiability or linearity of the constraint functions and dynamics, which limits the generality of the PI-MPC formulation. Moreover, projection distorts the sample density, potentially biasing the importance weights; the resulting solution may therefore deviate from the true constrained optimum obtained by the VI-based approach described next.
3.3.2 Constrained Particle-based Variational Inference
A more principled alternative is to embed state constraints directly into the variational inference problem, yielding a constrained KL minimization over Eq. (9). This line of work builds on constrained VI methods from the machine learning literature [liu2021sampling, zhang2022sampling], often implemented via SVGD [liu2016stein], which represents the target distribution as particles and updates them along the gradient of the KL divergence.
Power et al.  [power2024constrained] extended the constrained SVGD framework of [zhang2022sampling] to handle multiple linear and nonlinear constraints, applying it to robotic planning problems by projecting particle updates onto the tangent space of the constraint manifold. Tabor et al.  [tabor2025constrained] proposed a more unified formulation that folds constraints into the SVGD kernel, avoiding explicit tangent-space projection.
Constrained VI is theoretically appealing: it directly approximates the optimal control distribution conditioned on constraint satisfaction and can, in principle, handle constraints on distributional statistics (e.g., expected constraint satisfaction). In practice, however, iterative particle updates incur significant computational overhead, and the additional hyperparameters (e.g., step size, constraint penalty weights) make tuning more delicate than for standard MPPI.
3.4 Applications to High-Dimensional Robots
Early PI-MPC studies focused on low-dimensional systems such as ground vehicles with simple action spaces. As the field moves toward more capable platforms. e.g., humanoids with 30+ joints, multi-fingered hands, and soft robots, the dimensionality of the control problem grows rapidly, and naive Monte Carlo sampling becomes prohibitively expensive. Scaling PI-MPC to high degrees of freedom (DoF) requires addressing two coupled challenges: the exponential growth of sample complexity with action-space dimension, and the difficulty of obtaining accurate dynamics models for systems with complex contact and deformation. This section reviews how recent work tackles these challenges for platforms including 4-Wheel Independent Drive and Steering (4WIDS) vehicles [aoki2024switching], manipulators [bhardwaj2022storm], quadruped robots [keshavarz2025control], and humanoid robots [xue2024full].
3.4.1 Sample Complexity in High DoF Robots
The curse of dimensionality is the central obstacle: as the action-space dimension grows, the volume of the sampling region expands exponentially, and the fraction of samples that land near an optimal mode shrinks correspondingly. Standard MPPI with joint-angle or joint-torque actions often fails to find good solutions within a feasible sample budget for systems.
The most common countermeasure is dimensionality reduction. Defining the action space in terms of spline control points, rather than per-timestep joint commands, reduces the effective dimensionality and has enabled PI-MPC on various humanoid platforms [howell2022predictive, pezzato2025sampling, miura2024spline]. Combining Halton sequences (for low-discrepancy coverage) with spline interpolation further improves sample efficiency [bhardwaj2022storm, pezzato2025sampling], though such smoothing techniques inevitably sacrifice some optimality by restricting the space of representable trajectories.
An alternative to dimensionality reduction is iterative refinement. Diffusion-style annealing [pan2024model] and adaptive importance sampling (AIS) [asmar2023model] sharpen the sample distribution over multiple passes, concentrating samples in high-quality regions. These methods have shown promising results, including real-world deployment on quadruped [keshavarz2025control] and humanoid [xue2024full] robots, but the additional iterations increase computational cost.
3.4.2 Dynamics Modeling for High DoF Robots
Accurate dynamics models are essential for PI-MPC, yet analytical modeling becomes intractable for high-DoF robots, especially those involving contact (e.g., legged locomotion, manipulation). A practical solution is to use GPU-accelerated physics simulators, e.g., MuJoCo [todorov2012mujoco], Isaac Lab [mittal2025isaaclab], and Brax [freeman1brax], for forward rollouts during planning [howell2022predictive, pezzato2025sampling, xue2024full]. Isaac Lab and Brax are particularly well-suited to PI-MPC because they support massively parallel simulation on GPUs, and their randomizable environments enable domain randomization for robustness.
An alternative is to learn the dynamics model from data and use it for planning, i.e., the model-based RL. Methods such as TD-MPC [hansen2022temporal, hansentd] learn compact latent-space dynamics and plan with MPPI, achieving strong performance even on high-DoF continuous-control tasks.
3.5 Hardware Acceleration
Real-time PI-MPC demands that all computation, i.e., sampling, rollout, cost evaluation, and weight update, complete within a few milliseconds. Because the sampling and rollout steps are embarrassingly parallel, the primary lever for meeting this budget is hardware-level parallelism. The choice of platform depends on the deployment context: CPUs suffice for low-dimensional systems with moderate sample counts, GPUs are the current workhorse for high-dimensional problems, and emerging architectures (e.g., FPGAs, quantum-inspired hardware) target ultra-low-latency or energy-constrained settings.
On the CPU side, the ROS 2 Navigation stack includes an MPPI-based local planner [macenski2020marathon2] implemented with the Eigen library, offering efficient single-threaded and SIMD-parallel execution suitable for embedded deployment. On the GPU side, machine-learning frameworks such as PyTorch and JAX provide high-level abstractions that make GPU-accelerated MPPI straightforward to implement [bou2023torchrl], though Python-level overhead limits throughput. Native CUDA kernel implementations in C++ eliminate this overhead and deliver substantially higher performance [vlahov2024mppi].
Beyond conventional CPUs and GPUs, recent work has begun exploring unconventional accelerators. Werthen-Brabants and Simoens et al.  [werthen2025ising] proposed solving the MPPI optimization on Ising machines, a form of quantum-inspired computing. Tanguy-Legac et al.  [tanguy2026domain] investigated dedicated FPGA and ASIC designs tailored to the MPPI dataflow.
3.6 Theoretical Analysis of MPPI
Despite its empirical success, PI-MPC lacks the mature theoretical toolkit available for gradient-based MPC or LQR. We must often tune parameters by trial and error, without formal guidance on how close the resulting solution is to the true optimum or how many samples are sufficient. The two central questions are: (i) how large is the optimality gap between the MPPI solution and the solution of the original optimal control problem, and (ii) under what conditions does MPPI converge, and at what rate? Progress on these questions not only informs algorithm design but is also essential for certifying PI-MPC in safety-critical applications.
3.6.1 Optimality Gap from Original Optimal Control Problem
Because MPPI minimizes the KL objective in Eq. (5) rather than the original cost in Eq. (1), its solution can deviate from the true optimum. Williams et al.  [williams2018information] identified a special case in which this gap vanishes: when the cost separates into state-dependent and input-dependent terms and the input cost is a quadratic weighted by the prior covariance, the MPPI solution coincides exactly with the optimal solution of the original problem. Trevisan et al.  [trevisan2024biased] generalized this result, showing that for any variational family, adding an appropriate bias correction to the cost function eliminates the gap.
More broadly, the optimality gap can be traced to the prior regularization term in Eq. (5). Appropriate prior design (Section 3.1) and temperature tuning can therefore reduce the gap in practice. Yi et al.  [yi2024covo] derived a closed-form covariance that maximizes the convergence rate, directly linking prior design to optimality. Homburger et al.  [homburger2025optimality] provided a more general analysis for deterministic problems, proving that as $\lambda\to 0$ the MPPI solution converges to the deterministic optimum and bounding the control-input and value-function errors as functions of $\lambda$.
3.6.2 Convergence Analysis
Yi et al.  [yi2024covo] proved that MPPI converges for quadratic costs in the infinite-sample limit, with the convergence rate depending on the temperature parameter and the prior covariance. They further suggested that contraction properties extend to strongly convex costs with nonlinear dynamics.
Because MPPI is a zeroth-order method, its convergence is intimately tied to sample complexity. Yoon et al.  [Yoon2022-jb] and Tao et al.  [tao2022path] derived probabilistic bounds on the estimation error of the optimal distribution, showing that smaller prior covariance and lower expected trajectory costs both reduce the number of samples needed. However, convergence depends jointly on sample count, temperature, prior covariance, and cost landscape in ways that are not yet fully disentangled.
3.7 Connection to Other Paradigms
PI-MPC does not exist in isolation; it shares deep structural connections with reinforcement learning, diffusion models, and game-theoretic control. Understanding these connections clarifies which tools from neighboring fields can be imported and where PI-MPC offers unique advantages. We highlight three principal threads: the Control-as-Inference bridge to RL, the emerging link between MPPI and diffusion processes, and applications to broader problem settings.
Control as Inference and Reinforcement Learning. The Control-as-Inference framework [levine2018reinforcement] provides a unified probabilistic perspective that encompasses both MPC and RL. Within this framework, RL algorithms such as Soft Actor-Critic (SAC) [haarnoja2018soft] and Maximum a Posteriori Policy Optimization (MPO) [abdolmaleki2018maximum] emerge from the same KL minimization that underlies PI-MPC; the key difference is that RL methods amortize the optimization over a learned policy, whereas PI-MPC solves it online at each time step. Model-based RL bridges the two: methods such as PETS [chua2018deep], Dreamer [hafner2019learning], and TD-MPC [hansen2022temporal, hansentd] learn dynamics and reward models while using PI-MPC planners (CEM or MPPI) for action selection. In this context, the stochastic nature of PI-MPC naturally provides the exploration variability that model-based RL requires.
Diffusion Models. MPPI and score-based diffusion processes [ho2020denoising] are closely related: both iteratively refine a noisy distribution toward a target. Pan et al.  [pan2024model] and Xue et al.  [xue2024full] showed that when an optimal control problem replaces training data as the target, MPPI computations can substitute for portions of the score function, enabling sample generation at each diffusion step without training. This diffusion-style annealing captures finer modal structure than single-pass MPPI and has demonstrated strong results on high-dimensional locomotion tasks. Li et al.  [li2025unifying] further attempt to unify MPPI, RL, and diffusion models under the lens of gradient ascent on smoothed energy with respect to Gibbs measures, offering a common theoretical foundation.
Other Applications. Beyond standard single-agent control, MPPI has been extended to a range of broader settings, including multi-agent and adversarial formulations. Streichenberg et al.  [streichenberg2023multi] study MPPI in multi-agent control, and Zhang et al.  [zhang2025zsg] reformulate stochastic zero-sum games as optimal control problems that can be solved with MPPI. MPPI has also been used for data augmentation in visual navigation [inglin2026less]. Taken together, these extensions suggest that the PI-MPC framework is broadly applicable whenever sequential decision-making can be cast as trajectory optimization.